# A localisation phase transition for the catalytic branching random walk

### v1

@@ 4.4 Checking Assumption (A’3)
Assume that there exists a finite set $L\subset E$ such that $\mathbb{P}(X(1)=y)>0$ for all $x,y\in L$, and such that the constant
$\eta_{2}=\inf\{\eta>0\colon\liminf_{t\uparrow\infty}\mathrm{e}^{-\eta t}\mathbb{P}_{x}(X(t)=x)\}$
is finite and does not depend on $x\in L$. If there exists $C>0$, $\eta_{1}>\eta_{2}$ and $\varphi\colon E\cup\{\partial\}\to[0,\infty)$ such that $\varphi_{|E}\geq 1$, $\varphi(\partial)=0$, $\sum_{y\in E\setminus\{x\}}q_{x,y}\varphi(y)<\infty$ for all $x\in E$, $\eta_{1}>\sup_{x\in E}q_{x,\partial}$ and
@@
[…]
We first prove that the assumptions of Theorem [CV23] are satisfied: Recall that, for all $x\in\mathbb{Z}^{d}$,
@@
=$=\begin{cases}(\lambda_{0}-1)\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] (\lambda-1)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}&\text{ if }\|x\|_{1}=1\\[5.0pt] (\lambda-1)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]&\text{ if }\|x\|_{1}>1.\end{cases}$
Instead of considering the jump process $X$ of generator $(Q_{x}-\delta_{x}+(1-Q_{x}(\mathbb{Z}^{d}))\delta_{\partial})_{x\in\mathbb{Z}^{d}}$, we look at the jump process $Y$ of generator $(\kappa Q_{x}-\kappa\delta_{x}+\kappa(1-Q_{x}(\mathbb{Z}^{d}))\delta_{\partial})_{x\in\mathbb{Z}^{d}}$. One can couple $X$ and $Y$ so that $(X(\kappa t))_{t\geq 0}=(Y(t))_{t\geq 0}$. In particular, (A3) holds for $X$ if and only if it holds for $Y$.
Now note that, for all $x\in E$,
@@
=$=\begin{cases}(\lambda_{0}-1-\kappa)\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] (\lambda-1-\kappa)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}=1\\[5.0pt] (\lambda-1-\kappa)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}>1.\end{cases}$
=$=\begin{cases}-\frac{1+\lambda}{1+\lambda_{0}}\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] -\big{(}\lambda_{0}-\lambda+\frac{1+\lambda}{1+\lambda_{0}}\big{)}\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}=1\\[5.0pt] -\big{(}\lambda_{0}-\lambda+\frac{1+\lambda}{1+\lambda_{0}}\big{)}\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}>1.\end{cases}$
@@
  • When the particle sits at $x$ and $\|x\|_{1}=1$, it jumps to a uniformly-chosen neighbouring site at rate $1$, it jumps to 00 with an additional rate of $\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}$, and it dies at rate $(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}$.
@@
Note in particular that the dying rate is the largest when $\|x\|_{1}>1$ and the lowest at $x=0$ (in fact, the particle does not dies when it sits at the origin). Also note the additional drift towards zero when the particle sits at a neighbouring site of the origin.
To prove (A3), we aim at applying Theorem 4.4. To do this, we need to check that $(Y(t))_{t\geq 0}$ satisfies Assumptions (F0-3). We fix
@@
which is equivalent to $2d-1+2d\lambda<\lambda_{0}$, which is our assumption. We choose $\varepsilon>0$ small enough so that $\rho_{2}<\rho_{1}-\varepsilon$. We then choose $K>\nicefrac{{1}}{{\varepsilon}}$ large enough so that, for all $x\geq K$,
$\Big{(}1+\frac{1}{x}\Big{)}^{\nicefrac{{3}}{{4}}}\leq 1+\frac{1}{x}.$ (4.16)
We now let $L=\{x\in\mathbb{Z}^{d}\colon\|x\|_{1}\leq K\}$ and $\varphi(x)=\max(1,\|x\|_{1}^{\nicefrac{{3}}{{4}}})$.
@@
$\liminf_{t\uparrow\infty}\mathrm{e}^{-\eta t}\mathbb{P}_{x}(X(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{2\eta}\liminf_{t\uparrow\infty}\mathrm{e}^{-\eta t}\mathbb{P}_{y}(X(t)=y),$
which implies that $\eta_{0}$ does not depend on $x\in L$; it only remains to show that it is finite. Now note that, if we let $L_{0}=\{\|x\|\leq 1\}$ for all $x\in L$, $t\geq 2$,
@@
$\mathbb{P}_{x}(Y(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{\rho_{2}(t-2)},$
which implies that $\liminf_{t\uparrow\infty}\mathbb{P}_{x}(Y(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{2\rho_{2}}>0$. Thus,
$\eta_{1}\leq\rho_{2}<\infty.$
Now, because the death rate outside of $L$ equals $\rho_{1}$ (see (4.14) for the definition of $\rho_{1}$), we get that, for all $x\notin L$, for all $t\geq 0$,
@@
where we have set $C=\max_{x\in L}\mathcal{L}\varphi(x)<\infty$ and $\eta_{1}=\rho_{1}-\varepsilon$. Recall that we have chosen $\varepsilon$ such that $\eta_{2}=\rho_{2}<\rho_{1}-\varepsilon=\eta_{1}$. We have thus shown that the assumptions of Theorem 4.4 hold for $L=\{\|x\|\leq K\}$ for some $K$ large enough and $\varphi_{1}=\|\cdot\|^{\nicefrac{{3}}{{4}}}=V^{\nicefrac{{1}}{{q}}}$ (for the choices of $V$ and $q$ made in Section 4.3). We thus get that $X$ admits a quasi-stationary distribution $\nu$ as required in (A’3). Furthermore, for any $\mu\in\mathcal{P}_{C}(E)$ such that $\|\mu-\nu\|_{TV}\leq\nu(L)/2$, we get that
[…]
and thus (A’3) holds.
Acknowledgements: We are very grateful to Denis Villemonais for guiding us through [CV23] and to Daniel Kious for early discussions on this project.

### v2

@@ 5.4 Checking Assumption (A’3)
Assume that there exists a finite set $L\subset E$ such that $\mathbb{P}_{x}(X(1)=y)>0$ for all $x,y\in L$, and such that the constant
$\eta_{2}=\inf\{\eta>0\colon\liminf_{t\uparrow\infty}\mathrm{e}^{\eta t}\mathbb{P}_{x}(X(t)=x)>0\}$
is finite and does not depend on $x\in L$. If there exists $C>0$, $\eta_{1}>\eta_{2}$ and $\varphi\colon E\cup\{\partial\}\to[0,\infty)$ such that $\varphi_{|E}\geq 1$, $\varphi(\partial)=0$, $\sum_{y\in E\setminus\{x\}}q_{x,y}\varphi(y)<\infty$ for all $x\in E$, and
@@
then the process $(X(t))_{t\geq 0}$ admits a quasi-stationary distribution $\nu_{\mathit{QSD}}$, which is the unique one satisfying $\nu_{\mathit{QSD}}\varphi<\infty$ and $\mathbb{P}_{\nu_{\mathit{QSD}}}(X(t)\in L)>0$ for some $t\geq 0$. Moreover, there exist constants $\alpha\in(0,1)$ and $K>0$ such that, for all probability measures $\mu$ on $E$ such that $\int\varphi\,\mathrm{d}\mu<\infty$ and $\mu(L)>0$,
$\|\mathbb{P}_{\mu}(X(t)\in\cdot|X(t)\neq\partial)-\nu_{\mathit{QSD}}\|_{\mathit{TV}}\leq K\alpha^{t}\cdot\frac{\int\varphi\,\mathrm{d}\mu}{\mu(L)}.$ (5.19)
Furthermore, if we set $\varpi_{t}(x)=\mathrm{e}^{\eta_{2}t}\mathbb{P}_{x}(X(t)\neq\partial)$, then there exists a function $\varpi:\mathbb{E}\mapsto[0,\infty)$ such that
$\|\varpi_{t}-\varpi\|_{L^{\infty}(\varphi)}\to 0,$
where
$L^{\infty}(\varphi)=\{f:E\mapsto R \colon\|f\|_{L^{\infty}(\varphi)}:=\sup_{x\in E}|f(x)|/\varphi(x)<\infty\}.$
Remark.
Equation (5.19) follows from taking $n_{0}=0$ in the definition of $\psi_{2}$ in [CV23, Theorem 5.1, see also Theorem 3.1]; a private communication by Denis Villemonais confirmed to us that “for all large enough $n_{0}$” in the conclusion of [CV23, Theorem 3.1] should be replaced by “for any $n_{0}$”. In particular, the choice of $n_{0}=0$ is valid.
We first prove that the assumptions of Theorem 5.4 are satisfied: Recall from (5.4) that, for all $x\in\mathbb{Z}^{d}$,
@@
$=\begin{cases}(\lambda_{0}-1)\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] (\lambda-1)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}&\text{ if }\|x\|_{1}=1\\[5.0pt] (\lambda-1)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]&\text{ if }\|x\|_{1}>1.\end{cases}$
Instead of considering the jump process $X$ of generator $(Q_{x}-\delta_{x}+(1-Q_{x}(\mathbb{Z}^{d}))\delta_{\partial})_{x\in\mathbb{Z}^{d}}$, we look at the jump process $Y$ of generator $(\kappa Q_{x}-\kappa\delta_{x}+\kappa(1-Q_{x}(\mathbb{Z}^{d}))\delta_{\partial})_{x\in\mathbb{Z}^{d}}$. One can couple $X$ and $Y$ so that $(X(\kappa t))_{t\geq 0}=(Y(t))_{t\geq 0}$.
Using the definition of $\kappa$ (see (5.1)) and (5.5), we get that, for all $x\in\mathbb{Z}^{d}$,
@@
$=\begin{cases}(\lambda_{0}-1-\kappa)\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] (\lambda-1-\kappa)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}=1\\[5.0pt] (\lambda-1-\kappa)\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}>1.\end{cases}$
$=\begin{cases}-\frac{1+\lambda}{1+\lambda_{0}}\delta_{0}+\frac{1+\lambda}{1+\lambda_{0}}\mathbb{E}[\delta_{x+\Delta}]&\text{ if }x=0\\[5.0pt] -\big{(}\lambda_{0}-\lambda+\frac{1+\lambda}{1+\lambda_{0}}\big{)}\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}\delta_{0}+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}=1\\[5.0pt] -\big{(}\lambda_{0}-\lambda+\frac{1+\lambda}{1+\lambda_{0}}\big{)}\delta_{x}+\mathbb{E}[\delta_{x+\Delta}]+(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}\big{)}\delta_{\partial}&\text{ if }\|x\|_{1}>1.\end{cases}$
@@
  • When the particle sits at $x$ and $\|x\|_{1}=1$, it jumps to a uniformly-chosen neighbouring site at rate $1$, it jumps to 0 with an additional rate of $\frac{1}{2d}\frac{\lambda_{0}-\lambda}{1+\lambda}$, and it dies at rate $(\lambda_{0}-\lambda)\big{(}1-\frac{1}{1+\lambda_{0}}-\frac{1}{2d}\frac{1}{1+\lambda}\big{)}$.
@@
Note in particular that the dying rate is the largest when $\|x\|_{1}>1$ and the lowest at $x=0$ (in fact, the particle does not die when it sits at the origin). Also note the additional drift towards zero when the particle sits at a neighbouring site of the origin.
To prove (A’3), we aim at applying Theorem 5.4 and thus check that $(Y(t))_{t\geq 0}$ satisfies its assumptions. We fix
@@
which is equivalent to $2d-1+2d\lambda<\lambda_{0}$, which is our assumption. We choose $\varepsilon>0$ small enough so that $\rho_{2}<\rho_{1}-\varepsilon$. We then choose $K>\nicefrac{{1}}{{\varepsilon}}$ and let $L=\{x\in\mathbb{Z}^{d}\colon\|x\|_{1}\leq K\}$. We also let $\varphi(x)=\max(1,\|x\|_{1}^{\nicefrac{{3}}{{4}}})$, for all $x\in\mathbb{Z}^{d}$ (and $\varphi(\partial)=0$, as required).
@@
$\liminf_{t\uparrow\infty}\mathrm{e}^{\eta t}\mathbb{P}_{x}(X(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{-2\eta}\liminf_{t\uparrow\infty}\mathrm{e}^{\eta t}\mathbb{P}_{y}(X(t)=y),$
which implies that $\eta_{2}$ does not depend on $x\in L$; it only remains to show that it is finite. Now note that, if we let $L_{0}=\{\|x\|\leq 1\}$ then, for all $x\in L$, $t\geq 2$,
@@
$\mathbb{P}_{x}(Y(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{-\rho_{2}(t-2)},$
which implies that $\liminf_{t\uparrow\infty}\mathrm{e}^{\rho_{2}t}\mathbb{P}_{x}(Y(t)=x)\geq\mathfrak{m}^{2}\mathrm{e}^{2\rho_{2}}>0$. Thus,
$\eta_{2}\leq\rho_{2}<\infty.$
We now let $\mathcal{L}$ be the infinitesimal generator of $(Y(t))_{t\geq 0}$. Recall that $\varphi(x)=\max(1,\|x\|_{1})$, for all $x\in\mathbb{Z}^{d}$. From now on, we write $\|\cdot\|$ in lieu of $\|\cdot\|_{1}$. Because the death rate outside of $L$ equals $\rho_{1}$ (see (5.20) for the definition of $\rho_{1}$), we get that, for all $x\notin L$, for all $t\geq 0$,
@@
where we have set $C=\max_{x\in L}\mathcal{L}\varphi(x)<\infty$ and $\eta_{1}=\rho_{1}-\varepsilon$. Recall that we have chosen $\varepsilon$ such that $\eta_{2}\leq\rho_{2}<\rho_{1}-\varepsilon=\eta_{1}$. We have thus shown that the assumptions of Theorem 5.4 hold for $L=\{\|x\|\leq K\}$ for some $K$ large enough and $\varphi=\max(1,\|\cdot\|^{\nicefrac{{3}}{{4}}})=V^{\nicefrac{{1}}{{q}}}$ (for the choices of $V$ and $q$ made in Section 5.3). We thus get that $Y$ admits a quasi-stationary distribution $\nu$ and that, for any $\mu\in\mathcal{P}_{C}(\mathbb{Z}^{d})$ such that $\|\mu-\nu\|_{TV}\leq\nu(L)/2$,
$\|\mathbb{P}_{\mu}(Y(t)\in\cdot|Y(t)\neq\partial)-\nu\|_{\mathit{TV}}\leq C\alpha^{t}\cdot\frac{\int\varphi\,\mathrm{d}\mu}{\mu(L)}\leq C\alpha^{t}\cdot\frac{C}{\nu(L)/2}.$
Because $X(t)=Y(\nicefrac{{t}}{{\kappa}})$, we get that
$\|\mathbb{P}_{\mu}(X(t)\in\cdot|X(t)\neq\partial)-\nu\|_{\mathit{TV}}=\|\mathbb{P}_{\mu}(Y(\nicefrac{{t}}{{\kappa}})\in\cdot|Y(\nicefrac{{t}}{{\kappa}})\neq\partial)-\nu\|_{\mathit{TV}}\leq C\alpha^{t/\kappa}\cdot\frac{C}{\nu(L)/2}.$
Thus, to prove that (A’3) holds, it only remains to prove that, for all $\alpha\in\mathcal{P}_{C}(\mathbb{Z}^{d})$, $\mathbb{P}_{\alpha}(X(t)\neq\partial)\geq\mathrm{e}^{-(1-c_{0})t}$, i.e. that
$\mathbb{P}_{\alpha}(Y(t)\neq\partial)\geq\mathrm{e}^{-\kappa(1-c_{0})t},$ (5.23)
for some constant $c_{0}>\theta=\nicefrac{{\lambda}}{{\kappa}}$. If we set $\varpi_{t}(x)=\mathrm{e}^{\eta_{2}t}\mathbb{P}_{x}(Y(t)\neq\partial)$ for all $t\geq 0$ and $x\in\mathbb{Z}^{d}$, then
$\mathrm{e}^{\eta_{2}t}\mathbb{P}_{\alpha}(Y(t)\neq\partial)=\alpha\cdot\varpi\bigg{(}=\int_{\mathbb{Z}^{d}}\varpi_{t}\mathrm{d}\alpha\bigg{)}.$
We apply the dominated convergence theorem to this integral: To do so, first note that, by Theorem 5.4, there exists a function $\varpi$ such that $\|\varpi_{t}-\varpi\|_{L^{\infty}(\varphi)}\to 0$ as $t\uparrow\infty$. This implies that, for all $\varepsilon>0$, there exists $t_{\varepsilon}$ such that, for all $t\geq t_{\varepsilon}$, for all $x\in\mathbb{Z}^{d}$,
$|\varpi_{t}(x)-\varpi(x)|\leq\varepsilon\varphi(x).$ (5.24)
Thus
$\big{|}\alpha\cdot(\varpi_{t}-\varpi)\big{|}\leq\varepsilon\alpha\cdot\varphi\leq C\varepsilon,$
because $\alpha\in\mathcal{P}_{C}(\mathbb{Z}^{d})$. This implies
$0\leq\alpha\cdot\varpi_{t}=\alpha\cdot\varpi+\alpha\cdot(\varpi_{t}-\varpi)\leq\|\varphi\|_{L^{\infty}(\varphi)}\alpha\cdot\varphi+C\varepsilon<\infty,$
because $\varpi\in L^{\infty}(\varphi)$ (as a limit of functions in $L^{\infty}(\varphi)$) and $\alpha\in\mathcal{P}_{C}(\varphi)$. Thus, by dominated convergence, we get that
$\mathrm{e}^{\eta_{2}t}\mathbb{P}_{\alpha}(Y(t)\neq\partial)=\alpha\cdot\varpi_{t}\to\alpha\cdot\varpi.$ (5.25)
We now prove that $\alpha\cdot\varpi>0$; this holds because $\varpi(x)>0$ for all $x\in\mathbb{Z}^{d}$. Indeed, reasoning as in (5.22), we have that, for all $x\in\mathbb{Z}^{d}$, for all $t>2$,
$\mathbb{P}_{x}(Y(t)\neq\partial)$ $\geq\mathbb{P}_{x}(Y(t)=x)\geq\sum_{z\in L_{0}}\mathbb{P}_{x}(Y(1)=0)\mathbb{P}_{0}(Y(t-2)=z)\mathbb{P}_{z}(Y(1)=x)$
$\geq\mathbb{P}_{x}(Y(1)=0)\big{(}\min_{z\in L_{0}}\mathbb{P}_{z}(Y(1)=x)\big{)}\mathbb{P}_{0}(Y(s)\in L_{0}\text{ for all }0\leq s\leq t-2)$
$\geq\mathbb{P}_{x}(Y(1)=0)\big{(}\min_{z\in L_{0}}\mathbb{P}_{z}(Y(1)=x)\big{)}\mathrm{e}^{-\rho_{2}(t-2)}.$
Because $\rho_{2}=\eta_{2}$,
$\varpi(x)=\lim_{t\uparrow\infty}\mathrm{e}^{\eta_{2}t}\mathbb{P}_{x}(Y(t)\neq\partial)\geq\mathbb{P}_{x}(Y(1)=0)\big{(}\min_{z\in L_{0}}\mathbb{P}_{z}(Y(1)=x)\big{)}\mathrm{e}^{2\eta_{2}}>0.$
This concludes the proof that $\alpha\cdot\varpi>0$. Thus, (5.25) implies that there exists $\varepsilon>0$ such that, for all $t\geq 0$,
$\mathrm{e}^{\eta_{2}t}\mathbb{P}_{\alpha}(Y(t)\neq\partial)>\varepsilon.$
Thus, to prove (5.23), it is enough to show that $\rho_{2}=\eta_{2}<\kappa(1-\theta)=\kappa-\lambda$. By (5.1),
$\kappa-\lambda=\frac{\lambda_{0}(\lambda_{0}-\lambda)}{1+\lambda_{0}}.$
Also, by (5.21),
$\eta_{2}=\frac{\lambda_{0}(\lambda_{0}-\lambda)}{1+\lambda_{0}}+\frac{2d-1}{2d}-\frac{\lambda_{0}-\lambda}{2d(1+\lambda)}.$
Thus $\eta_{2}<\kappa-\lambda$ if and only if
$\frac{2d-1}{2d}<\frac{\lambda_{0}-\lambda}{2d(1+\lambda)}\quad\Leftrightarrow\quad\lambda_{0}>2d-1+2d\lambda,$
which holds, by assumption.
Acknowledgements:
We are very grateful to Denis Villemonais for guiding us through [CV23], Daniel Kious for early discussions on this project, Marcel Ortgiese for comments on the discussion of the literature, and Ofer Zeitouni for insightful comments, which led to Proposition 2.5.